OpenStudy (anonymous):

the area of a triangle is 338 square yards. the height of the triangle is four times its base. find the height and base of the triangle.

OpenStudy (anonymous):

A= 1/2 B*H 338= 1/2 B*4B 338=1/2(4B^2) 676= 4B^2 169=B^2 B=13 H= 4*13 H=52